The management of Advanced Alternative Power Inc. is considering two capital investment projects. The estimated net cash flows from each project are as follows:

Year Wind Turbines Biofuel Equipment
1 $160,000 $340,000
2 160,000 340,000
3 160,000 340,000
4 160,000 340,000

The wind turbines require an investment of $485,920, while the biofuel equipment requires an investment of $970,700. No residual value is expected from either project.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

Required:

1a. Compute the net present value for each project. Use a rate of 10% and the present value of an annuity of $1 in the table above. If required, use the minus sign to indicate a negative net present value. If required, round to the nearest whole dollar.

Wind Turbines Biofuel Equipment
Present value of annual net cash flows $ $
Less amount to be invested
Net present value $ $

1b. Compute a present value index for each project. If required, round your answers to two decimal places.

Present Value Index
Wind Turbines
Biofuel Equipment

2. Determine the internal rate of return for each project by (a) computing a present value factor for an annuity of $1 and (b) using the present value of an annuity of $1 in the table above. If required, round your present value factor answers to three decimal places and internal rate of return to the nearest whole percent.

Wind Turbines Biofuel Equipment
Present value factor for an annuity of $1
Internal rate of return % %
Answer #1

1a. Compute the net present value for each project. Use a rate of 10% and the present value of an annuity of $1 in the table above. If required, use the minus sign to indicate a negative net present value. If required, round to the nearest whole dollar.

Wind Turbines Biofuel Equipment
Present value of annual net cash flows $160000*3.17 = 507200 $340000*3.17 = 1077800
Less amount to be invested -485920 -970700
Net present value $21280 $107100

1b. Compute a present value index for each project. If required, round your answers to two decimal places.

Present Value Index
Wind Turbines 507200/485920 = 1.04
Biofuel Equipment 1077800/970700 = 1.11

2. Determine the internal rate of return for each project by (a) computing a present value factor for an annuity of $1 and (b) using the present value of an annuity of $1 in the table above. If required, round your present value factor answers to three decimal places and internal rate of return to the nearest whole percent.

Wind Turbines Biofuel Equipment
Present value factor for an annuity of $1 485920/160000 = 3.037 970700/340000 = 2.855
Internal rate of return 12 % 15 %